# AGZO: Activation-Guided Zeroth-Order Optimization for LLM Fine-Tuning

### v1

@@ 4.1 AGZO Algorithm
Activation-informed subspace.
Given a target rank $r\ll\min\{d_{\text{out}},d_{\text{in}},m\}$, we approximate the top $r$ left singular vectors of $H_{\ell}$ via a few steps of power iteration on $H_{\ell}H_{\ell}^{\top}$, using only matrix–matrix products with $H_{\ell}$ and $H_{\ell}^{\top}$. The routine in Algorithm 1 takes the current activation matrix $H$ and returns an orthonormal basis $A\in\mathbb{R}^{d_{\text{in}}\times r}$ whose columns span a rank-$r$ subspace of $\mathrm{col}(H)$. Once $A_{\ell}$ is computed for the current minibatch, $H_{\ell}$ itself is discarded.
@@
Perturbations and zeroth order estimator.
Given $A_{\ell}\in\mathbb{R}^{d_{\text{in}}\times r_{\ell}}$, AGZO samples a low-rank perturbation for each linear layer by drawing a left factor $R_{\ell}\in\mathbb{R}^{d_{\text{out}}\times r_{\ell}}$ with i.i.d. standard normal entries and setting
@@
$\widehat{\nabla}_{W_{\ell}}f^{\text{AGZO}}(W;B)\;=\;\frac{f_{+}-f_{0}}{\mu}\,\Delta_{\ell},\qquad\ell=1,\dots,L.$ (13)
@@
17: Set projected gradient scaler $g\leftarrow(f_{+}-f_{0})/\mu$.
@@
In practice we use a very small number of power-iteration steps per layer ($K=3$), which adds only a few matrix multiplications on top of the forward pass. The dominant cost per iteration is thus forward evaluations, as in MeZO and LOZO.
4.2 Memory Usage
We briefly compare the ZO-specific memory overhead of MeZO, LOZO, and AGZO, ignoring the baseline cost of storing model parameters.
MeZO does not store perturbations and regenerates them from random seeds, incurring essentially no ZO-specific memory overhead. LOZO maintains both $U_{\ell}$ and $V_{\ell}$, with memory $\mathcal{O}((d_{\text{out}}+d_{\text{in}})r_{\ell})$ per layer. AGZO only stores the activation-informed bases $A_{\ell}$, with ZO-specific memory $\mathcal{O}(d_{\text{in}}r)$ per layer; activation matrices are discarded immediately after subspace extraction, and the left factors $R_{\ell}$ and perturbations $\Delta_{\ell}$ are used only within a single iteration.
In our experiments, we choose $r_{\ell}=1$, so the number of additional parameters in $A_{\ell}$ is negligible compared with the size of the full weight matrix $W_{\ell}$ (which has $d_{\text{out}}d_{\text{in}}$ parameters). Consequently, the extra memory required by AGZO is far smaller than the model memory itself.
5 Theoretical analysis of AGZO
We now analyze the AGZO estimator introduced in ection 4. The goal is to understand (i) its mathematical essence and how far it is from the true gradient, and (ii) how its directional quality compares to MeZO. We focus on linear layers where AGZO uses low-rank perturbations as in (10) and the zeroth order estimator in (16); layers perturbed with dense Gaussian noise behave as in standard two-point ZO and are omitted from the analysis.
5.1 Subspace-smoothing view of AGZO
@@
Condition on the subspace bases $A:=\{A_{\ell}\}$ computed at the current iterate $W$. Let $\Delta(W,R)$ be the random perturbation with layerwise blocks
$\Delta_{\ell}(W,R)=R_{\ell}A_{\ell}^{\top},\qquad R_{\ell}\sim\mathcal{N}(0,I_{m_{\ell}\times r_{\ell}})\text{ i.i.d.\ across }\ell,$
and define the subspace-smoothed objective
@@
The AGZO estimator for layer $\ell$ can be written as
where
$\phi(W,\Delta;B)=\frac{f(W+\mu\Delta,B)-f(W,B)}{\mu}.$ (16)
@@
Assume $F$ is differentiable and the usual interchange of gradient and expectation is valid. Conditioned on $A$, the AGZO estimator satisfies, for each linear layer $\ell$,
$\mathbb{E}_{R,B}\Big[\widehat{\nabla}^{\mathrm{AGZO}}_{W_{\ell}}(W;B)\,\Big|\,A\Big]=\nabla_{W_{\ell}}F_{\mu,A}(W)\,A_{\ell}A_{\ell}^{\top}.$ (17)
@@
Under standard smoothness assumptions, the difference between $\nabla F_{\mu,A}$ and $\nabla F$ after projection is small.
@@
In particular, the bias from smoothing vanishes linearly as $\mu\to 0$ once projected onto $\mathrm{span}(A_{\ell})$.
@@
The remaining component of the bias comes from projecting $\nabla_{W_{\ell}}F(W)$ onto $\mathrm{span}(A_{\ell})$. Recall the gradient factorization (8) and the subspace analsis (9). In AGZO, $A_{\ell}$ is constructed to approximate a low-rank activation subspace for $\mathrm{col}(H_{\ell})$. For exposition, consider the idealized case where this subspace exactly supports the gradient.
@@
5.2 Directional quality and comparison with MEZO
Since the update length can be tuned by step size, the effectiveness of a gradient estimator mainly depends on its directional quality: how well its direction aligns with the true gradient. This subsection analyzes the expected cosine similarity between the estimator and the true gradient, and compares AGZO with MeZO in an idealized noiseless setting.
Let $G\in\mathbb{R}^{m\times n}$ denote the true gradient $\nabla_{W_{\ell}}F(W)$, and let $\widehat{G}$ be the approximated gradient. Define the cosine similarity
$\cos(\widehat{G},G):=\frac{\langle\widehat{G},G\rangle_{F}}{\|\widehat{G}\|_{F}\,\|G\|_{F}},\qquad\langle A,B\rangle_{F}:=\mathrm{tr}(A^{\top}B).$ (21)
We analyze the expected cosine similarity in a noiseless setting ($\mu\to 0$) where the finite difference oracle returns exact directional derivatives and stochastic minibatch noise is ignored. In this regime $\widehat{G}$ becomes a linear function of the Gaussian direction, which allows a closed-form computation.
AGZO.
For AGZO, focus on a single layer and write $G\in\mathbb{R}^{m\times n}$ and $A\in\mathbb{R}^{n\times r}$ (with orthonormal columns) for simplicity. The low-rank perturbation is $\Delta=RA^{\top}$ with $R\sim\mathcal{N}(0,I_{m\times r})$, and the corresponding estimator in the limit $\mu\to 0$ has the form
@@ 5.2 Directional quality and comparison with MEZO
Theorem 5.4.
Let $G\in\mathbb{R}^{m\times n}$ be fixed and $A\in\mathbb{R}^{n\times r}$ have orthonormal columns. Let $\widehat{G}^{\mathrm{AGZO}}_{0}$ be the noiseless AGZO estimator constructed from $\Delta=RA^{\top}$ with $R\sim\mathcal{N}(0,I_{m\times r})$. Then
$\mathbb{E}_{R}\big[\cos\big(\widehat{G}^{\mathrm{AGZO}}_{0},G\big)\big]=\beta_{mr}\,\frac{\|GA\|_{F}}{\|G\|_{F}},$ (23)
@@
and $\beta_{D}$ decays on the order of $1/\sqrt{D}$ as $D$ increases.
@@
The factor $\|GA\|_{F}/\|G\|_{F}$ has a natural geometric interpretation: it is precisely the fraction of gradient Frobenius energy captured by the AGZO subspace, since $\|GA\|_{F}=\|GAA^{\top}\|_{F}$ (See remark in Appendix). Thus AGZO benefits both from working in a lower effective dimension $mr$ (through $\beta_{mr}$) and from aligning its perturbation subspace with directions where $G$ has large energy.
MEZO.
For the MeZO baseline with full Gaussian directions, the estimator has the same form but with $A=I_{n}$ and $\Delta=R\in\mathbb{R}^{m\times n}$ dense. Theorem 5.4 then yields the following corollary.
Corollary $5.$
In the same setting as above, consider the noiseless MeZO estimator $\widehat{G}^{\mathrm{MEZO}}_{0}$ constructed from dense Gaussian directions $\Delta=R$ with $R\sim\mathcal{N}(0,I_{m\times n})$. Then
$\mathbb{E}_{R}\big[\cos\big(\widehat{G}^{\mathrm{MEZO}}_{0},G\big)\big]=\beta_{mn},$ (26)
where $\beta_{D}$ is as in (24). This corresponds to the special case of Proposition 5.4 with $A=I_{n}$ and $\|GA\|_{F}/\|G\|_{F}=1$.
Comparison under activation spectral concentration.
To compare AGZO and MeZO more explicitly, we recall facts about $\beta_{D}$ (proved in the appendix).
Lemma 5.6.
$\sqrt{\frac{2}{\pi D}}\;\leq\;\beta_{D}\;\leq\;\sqrt{\frac{2}{\pi(D-1)}}.$ (27)
In particular, $\beta_{mr}\approx c/\sqrt{mr}$ and $\beta_{mn}\approx c/\sqrt{mn}$ for a constant $c\approx\sqrt{2/\pi}$, so working in a much smaller subspace (with $r\ll n$) already improves the expected cosine similarity by a factor around $\sqrt{n/r}$.
Proof.
See Appendix A.2.1. ∎
The second ingredient is the energy ratio $\|GA\|_{F}/\|G\|_{F}$. In the noiseless case, Theorem 5.4 and Corollary 5.5 give that AGZO has larger expected cosine similarity than MeZO if and only if
$\frac{\|GA\|_{F}}{\|G\|_{F}}\;>\;\frac{\beta_{mn}}{\beta_{mr}}.$ (28)
By Lemma 5.6, the right-hand side admits the bound
$\frac{\beta_{mn}}{\beta_{mr}}\;<\;\frac{\sqrt{mr}}{\sqrt{mn-1}},$ (29)
so a sufficient condition for AGZO to outperform MeZO is
$\left(\frac{\|GA\|_{F}}{\|G\|_{F}}\right)^{2}\;>\;\frac{r}{n-1/m}.$ (30)
For a given layer, let $H_{\ell}\in\mathbb{R}^{n_{\ell}\times b}$ be the activation matrix and
$H_{\ell}=U_{\ell}\Sigma_{\ell}V_{\ell}^{\top},\qquad\Sigma_{\ell}=\mathrm{diag}(\sigma_{\ell,1},\dots,\sigma_{\ell,s_{\ell}}),$
its compact SVD with rank $s_{\ell}$ and singular values $\sigma_{\ell,1}\geq\dots\geq\sigma_{\ell,s_{\ell}}>0$. The layer gradient admits the factorization $\nabla_{W_{\ell}}F(W)=Q_{\ell}H_{\ell}^{\top}$, and we set
@@
AGZO chooses $A_{\ell}$ as the leading $r_{\ell}$ left singular vectors, $A_{\ell}=U_{\ell}^{(r_{\ell})}$. A direct calculation then yields
$\frac{\|G_{\ell}A_{\ell}\|_{F}^{2}}{\|G_{\ell}\|_{F}^{2}}=\frac{\sum_{i=1}^{r_{\ell}}B_{\ell,ii}\,\sigma_{\ell,i}^{2}}{\sum_{i=1}^{s_{\ell}}B_{\ell,ii}\,\sigma_{\ell,i}^{2}},$ (31)
so (30) is equivalent to
$\frac{\sum_{i=1}^{r_{\ell}}B_{\ell,ii}\,\sigma_{\ell,i}^{2}}{\sum_{i=1}^{s_{\ell}}B_{\ell,ii}\,\sigma_{\ell,i}^{2}}\;>\;\frac{r_{\ell}}{n_{\ell}-1/m_{\ell}}.$ (32)
This expression makes precise the intuition that AGZO benefits when the gradient energy is concentrated in the leading singular irections of the activation matrix.
The following theorem summarizes the consequence of the spectral assumptions we use in the analysis.
@@
Consider a layer with gradient factorization $\nabla_{W_{\ell}}F(W)=Q_{\ell}H_{\ell}^{\top}$ and compact SVD $H_{\ell}=U_{\ell}\Sigma_{\ell}V_{\ell}^{\top}$ of rank $s_{\ell}<n_{\ell}$ with singular values $\{\sigma_{\ell,i}\}_{i=1}^{s_{\ell}}$. Let $A_{\ell}=U_{\ell}^{(r_{\ell})}$ be the AGZO subspace and $B_{\ell}=V_{\ell}^{\top}Q_{\ell}^{\top}Q_{\ell}V_{\ell}$. If the average of its first $r$ diagonal entries is not less than the average of all diagonal entries, i.e.,
$\frac{1}{r_{\ell}}\sum_{i=1}^{r_{\ell}}B_{\ell,ii}\;\geq\;\frac{1}{s_{\ell}}\sum_{i=1}^{s_{\ell}}B_{\ell,ii},$ (33)
and that $H_{\ell}$ is low-rank, i.e., $s_{\ell}<n_{\ell}$. Then
$\frac{\sum_{i=1}^{r_{\ell}}B_{\ell,ii}\,\sigma_{\ell,i}^{2}}{\sum_{i=1}^{s_{\ell}}B_{\ell,ii}\,\sigma_{\ell,i}^{2}}\;>\;\frac{r_{\ell}}{s_{\ell}}\;\geq\;\frac{r_{\ell}}{n_{\ell}-1/m_{\ell}},$ (34)
and hence, in the noiseless single-query setting,
@@
Theorem 5.7 formalize the intuition behind AGZO: for layers whose gradients concentrate in low-rank activation subspaces, AGZO produces gradient estimates that are better aligned in direction than those of a dense Gaussian ZO method such as MeZO, while requiring only low-rank perturbations.
@@ 6 Experiments
Table 1: Experiments on Qwen3-0.6b. The best results are shown in bold except for FO.
@@
Table 2: Experiments on Pangu-1B. The best results are shown in bold except for FO.
@@
Avg. 0.751 0.720 0.706 0.687 0.595 0.640
This section evaluates AGZO on instruction-tuned language models under memory constraints. The experiments cover multiple tasks, including the SuperGLUE benchmark (Wang et al., 2019) and other datasets. We consider two model scales, Qwen3-0.6B and Pangu-1B, and compare against established zeroth-order baselines (MeZO, LoZO), as well as non-training baselines (zero-shot prompting and in-context learning, denoted as ICL). When feasible, we also include a first-order fine-tuning baseline (FO) using standard backpropagation. For ZO optimizer, we finetune the LLM for 20000 steps, and for FO optimizer, we use 1000 steps.
All experiments are run on an Ubuntu machine equipped with two NVIDIA RTX 3090 GPUs. All ZO methods update the full set of trainable parameters and use the same data preprocessing and evaluation pipeline. For fair comparison, each method is tuned via grid search over its key hyperparameters, and we report the best configuration selected on the validation set.

### v2

@@ 4.1 AGZO Algorithm
Activation-informed subspace. Given a target rank $r\ll\min\{d_{\text{out}},d_{\text{in}},m\}$, we approximate the top $r$ left singular vectors of $H_{\ell}$ via a few steps of power iteration on $H_{\ell}H_{\ell}^{\top}$, using only matrix–matrix products with $H_{\ell}$ and $H_{\ell}^{\top}$ (Bentbib & Kanber, 2015). The routine in Algorithm 1 takes the current activation matrix $H$ and returns an orthonormal basis $A\in\mathbb{R}^{d_{\text{in}}\times r}$ whose columns span a rank-$r$ subspace of $\mathrm{col}(H)$. Once $A_{\ell}$ is computed for the current minibatch, we discard $H_{\ell}$ to reduce memory consumption.
@@
Perturbations and zeroth order estimator. Given $A_{\ell}\in\mathbb{R}^{d_{\text{in}}\times r_{\ell}}$, AGZO samples a low-rank perturbation for each linear layer by drawing a left factor $R_{\ell}\in\mathbb{R}^{d_{\text{out}}\times r_{\ell}}$ with i.i.d. standard normal entries and setting
@@
$\widehat{\nabla}_{W_{\ell}}f^{\text{AGZO}}(W;B)\;=\;\frac{f_{+}-f_{0}}{\mu}\,\Delta_{\ell},\quad\ell=1,\dots,L.$ (10)
@@
17: Set projected gradient scalar $g\leftarrow(f_{+}-f_{0})/\mu$.
@@
In practice, a small number of power-iteration steps ($K=3$) per layer is enough to get satisfactory approximation, which adds only a few matrix multiplications on top of the forward pass. The dominant cost per iteration is thus forward evaluations, as in MeZO and LOZO.
4.2 Memory Usage Analysis
We analyze the memory footprint of each method, focusing on the optimization overhead—defined as the storage required beyond the fixed model parameters. Both MeZO and LOZO incur essentially no overhead, as their perturbations (whether isotropic or low-rank factors) are generated from random seeds and can be regenerated on the fly. AGZO requires storing the activation-informed basis $A_{\ell}\in\mathbb{R}^{d_{\text{in}}\times r}$ for each layer, as it depends on the input data and cannot be recovered from a seed. However, this overhead is negligible compared to the model size: for a weight matrix $W_{\ell}$ with $d_{\text{out}}\times d_{\text{in}}$ parameters, the basis $A_{\ell}$ requires only $d_{\text{in}}\times r$. With $r\ll d_{\text{out}}$, this consumes a tiny fraction of the memory needed for the weights themselves.
In our experiments, we set $r=1$. This choice minimizes the storage overhead for the basis $A_{\ell}$ to the theoretical lower bound. More importantly, since the AGZO perturbation is stochastic within the subspace spanned by $A_{\ell}$, restricting the rank to 1 forces the random exploration to concentrate entirely on the single most dominant direction of the activation energy. This prevents the update signal from being diluted across less significant components.
5 Provable Superiority of AGZO
We now analyze the AGZO estimator introduced in section 4. The goal is to understand (i) its mathematical essence and how far it is from the true gradient, and (ii) how its directional quality compares to MeZO. We focus on linear layers where AGZO uses low-rank perturbations as in Eq. (7) and the zeroth order estimator in Eq. (10).
5.1 AGZO is A Projected Gradient Estimator
@@
Condition on the subspace bases $A:=\{A_{\ell}\}$ computed at the current iterate $W$. Let $\Delta(W,R)$ be the random perturbation defined in Eq. (7) and define the subspace-smoothed objective
@@
Assume $F$ has $L$-Lipschitz gradient. The AGZO estimator satisfies, for each linear layer $\ell$,
$\mathbb{E}_{R,B}\Big[\widehat{\nabla}^{\mathrm{AGZO}}_{W_{\ell}}(W;B)\Big]=\nabla_{W_{\ell}}F_{\mu,A}(W)\,A_{\ell}A_{\ell}^{\top}.$ (12)
@@
Under standard smoothness assumptions, the difference between $\nabla F_{\mu,A}$ and $\nabla F$ after projection vanishes linearly as $\mu\to 0$.
@@
The remaining component of the bias comes from projecting $\nabla_{W_{\ell}}F(W)$ onto $\mathrm{span}(A_{\ell})$. Recall the gradient factorization (5) and the subspace analysis (6). In AGZO, $A_{\ell}$ is constructed to approximate a low-rank activation subspace for $\mathrm{col}(H_{\ell})$. For exposition, consider the idealized case where this subspace exactly supports the gradient.
@@
5.2 AGZO is A High-Precision Gradient Estimator
Since the update length can be tuned by step size, the effectiveness of a gradient estimator mainly depends on its directional quality: how well its direction aligns with the true gradient. This subsection analyzes the expected cosine similarity between the estimator and the true gradient, and compares AGZO with MeZO in a noiseless setting.
Let $G\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ denote the true gradient $\nabla_{W_{\ell}}F(W,B)$, and let $\widehat{G}$ be the approximated gradient. Define the cosine similarity:
$\cos(\widehat{G},G):=\frac{\langle\widehat{G},G\rangle_{F}}{\|\widehat{G}\|_{F}\,\|G\|_{F}},\quad\langle\widehat{G},G\rangle_{F}:=\mathrm{tr}(\widehat{G}^{\top}G).$ (16)
We analyze the expected cosine similarity in a noiseless setting ($\mu\to 0$) where the finite difference oracle returns exact directional derivatives and stochastic minibatch noise is ignored.
Theorem 5.4.
Let $G\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ be fixed and $A\in\mathbb{R}^{d_{\text{in}}\times r}$ have orthonormal columns. Let $\widehat{G}^{\mathrm{AGZO}}_{0}$ be the noiseless AGZO estimator, which has the form:
@@ 5.2 AGZO is A High-Precision Gradient Estimator
Then
$\mathbb{E}_{R}\big[\cos\big(\widehat{G}^{\mathrm{AGZO}}_{0},G\big)\big]=\beta_{d_{\text{out}}r}\,\frac{\|GA\|_{F}}{\|G\|_{F}},$ (18)
@@
and for any $D\geq 2$, $\beta_{D}$ satisfies the tight bounds:
$\sqrt{\frac{2}{\pi D}}\;\leq\;\beta_{D}\;\leq\;\sqrt{\frac{2}{\pi(D-1)}}.$ (20)
@@
The factor $\|GA\|_{F}/\|G\|_{F}$ has a natural geometric interpretation: it is precisely the fraction of gradient Frobenius energy captured by the AGZO subspace, since $\|GA\|_{F}=\|GAA^{\top}\|_{F}$ (See remark A.6 in Appendix). Thus AGZO benefits both from working in a lower effective dimension $mr$ (through $\beta_{mr}$) and from aligning its perturbation subspace with directions where $G$ has large energy.
For the MeZO baseline with Gaussian perturbations, the estimator has the same form but with $A=I_{d_{\text{in}}}$ and $\Delta=R\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ dense. Theorem 5.4 then yields the following corollary.
Corollary 5.5.
Consider the noiseless MeZO estimator $\widehat{G}^{\mathrm{MEZO}}_{0}$ constructed from dense Gaussian directions $\Delta=R$ with $R\sim\mathcal{N}(0,I_{d_{\text{out}}\times d_{\text{in}}})$. Then
$\mathbb{E}_{R}\big[\cos\big(\widehat{G}^{\mathrm{MEZO}}_{0},G\big)\big]=\beta_{d_{\text{out}}d_{\text{in}}}.$ (21)
This corresponds to the special case of Theorem 5.4 with $A=I_{d_{\text{in}}}$ and $\|GA\|_{F}/\|G\|_{F}=1$.
To compare AGZO and MeZO explicitly, we analyze their expected cosine similarity to the true gradient. Consider a layer with gradient factorization $\nabla_{W_{\ell}}F(W)=Q_{\ell}H_{\ell}^{\top}$. Let the compact SVD of the activation matrix be $H_{\ell}=U_{\ell}\Sigma_{\ell}V_{\ell}^{\top}$. We define the interaction matrix between the upstream gradient and activation subspaces as:
@@
This matrix captures the energy distribution of the gradient. Specifically, the diagonal entry $B_{\ell,ii}$ quantifies the energy of the upstream gradient projected onto the $i$-th principal component of the activation inputs.
We now state our main result. It demonstrates that unless the gradient signal is adversarially aligned with the smallest singular values of the activation, AGZO provably outperforms MeZO.
@@
Consider a layer where the activation matrix $H_{\ell}$ is low-rank (i.e., rank $s_{\ell}<d_{\text{in}}$). Assume the upstream gradient energy is broadly distributed such that the average energy along the top-$r_{\ell}$ directions is not less than the global average:
$\frac{1}{r_{\ell}}\sum_{i=1}^{r_{\ell}}B_{\ell,ii}\;\geq\;\frac{1}{s_{\ell}}\sum_{i=1}^{s_{\ell}}B_{\ell,ii}.$ (22)
Then, the AGZO provably yields a higher expected cosine similarity to the true gradient than MeZO :
@@
Furthermore, the performance gap widens as the activation singular values become more heterogeneous (i.e., faster decay).
@@
Theorem 5.6 formalizes the intuition behind AGZO: for layers where gradients concentrate in low-rank activation subspaces, AGZO produces update directions that are significantly better aligned with the true gradient than dense isotropic baselines.
@@ 6 Experiments
Table 1: Experiments on Qwen3-0.6b. Bold: ZO’s best results.
@@
Table 2: Experiments on Qwen3-4b.
Task AGZO MEZO LOZO Zero ICL
SST2 0.892 0.875 0.866 0.649 0.887
CB 0.875 0.857 0.857 0.375 0.821
BoolQ 0.820 0.823 0.822 0.790 0.827
MultiRC 0.853 0.850 0.852 0.765 0.849
RTE 0.848 0.837 0.801 0.805 0.835
WiC 0.678 0.666 0.659 0.595 0.615
Table 3: Experiments on Pangu-1B.
@@
This section evaluates AGZO on fine-tuning LLMs under practical memory constraints. The experiments cover multiple tasks, including the SuperGLUE benchmark (Wang et al., 2019) and other datasets. We conduct our evaluation on the Qwen3 family (0.6B and 4B scales) (Yang et al., 2025) and the openPangu  (Chen et al., 2025a) model. In particular, we select openPangu-embedded-1B (Rang et al., 2025) (denoted as Pangu-1B)  [^1] https://huggingface.co/FreedomIntelligence/openPangu-Embedded-1B, a model specifically designed for efficient inference on edge devices (we use the GPU variant in our experiments [^2] https://modelscope.cn/models/wangrongsheng/openPangu-Embedded-1B-model-GPU).
We compare AGZO against established zeroth-order baselines (MeZO, LOZO) as well as non-training baselines (zero-shot prompting and in-context learning, denoted as ICL). Additionally, we include a first-order fine-tuning baseline (FO) using standard backpropagation whenever memory constraints allow. ZO optimizers are fine-tuned for 20,000 steps, whereas the FO baseline is trained for 1,000 steps.
We build two testbeds with different computation platforms.
  1. 1.
    Testbed one is an Ubuntu machine equipped with two NVIDIA RTX 3090 GPUs.
  2. 2.
    Testbed two is an EulerOS machine equipped with eight Ascent 910B2 NPUs.
To ensure a fair comparison, all ZO methods update the full set of trainable parameters and share identical data preprocessing and evaluation pipelines. Regarding hyperparameters, we fix the smoothing parameter $\mu$ to $1\times 10^{-7}$ for all ZO methods on Qwen3 and set $\mu=1\times 10^{-4}$ for Pangu-1B, while the learning rate is determined via grid search based on validation set performance. Further experimental details are provided in Appendix B.